# An Oscillation-Free Real Fluid Quasi-Conservative Finite Volume Method for Transcritical and Phase-Change Flows

### v1

@@ Abstract
Abstract    A new Real Fluid Quasi-Conservative (RFQC) finite volume method is developed to address the numerical simulation of real fluids involving shock waves in transcritical and phase-change flows. To eliminate the spurious pressure oscillations inherent in fully conservative schemes, we extend the classic quasi-conservative method, originally designed for two-phase flows, to real fluids governed by arbitrary equations of state (EoS). The RFQC method locally linearizes the real fluid EoS at each grid point and time step, constructing and evolving the frozen Grüneisen coefficient $\Gamma$ and the linearization remainder $E_{0}$ via two advection equations. At the end of each time step, the evolved $\Gamma$ and $E_{0}$ are utilized to reconstruct the oscillation-free pressure field, followed by a thermodynamic re-projection applied to the conserved variables. Theoretical analysis demonstrates that, in smooth regions, the energy conservation error of the RFQC method is a high-order term relative to the spatial reconstruction truncation error. In discontinuous regions, this error is determined by the entropy increase rate, thereby maintaining consistency with the inherent truncation error of shock-capturing methods. A series of numerical tests verifies that the method can robustly simulate complex flow processes with only minor energy conservation errors, including transcritical flows, phase transitions, and shock-interface interactions. The RFQC method is proven to be both accurate and robust in capturing shock waves and phase transitions.
@@ 3.4 Re-Projection Error Analysis
In the thermodynamic re-projection step, only the total energy is modified. We define the resulting energy conservation error, $\epsilon_{p}=(\rho e_{t})^{n+1}_{temp}-(\rho e_{t})^{n+1}$, as the re-projection error. We now analyze the convergence properties of this error.
The energy $(\rho e_{t})^{n+1}_{temp}$ obtained by natural evolution of the conservation laws satisfies:
@@
$\epsilon_{p}=\rho^{n+1}\left(e_{real}(\rho^{n+1},p^{n+1})-\left(\frac{p^{n+1}(\xi)^{n+1}_{temp}+(E_{0})^{n+1}_{temp}}{\rho^{n+1}}\right)\right)$ (24)
We now analyze the temporal evolution of the re-projection error $\epsilon_{p}$. For clarity of derivation, superscripts and subscripts are omitted in the derivation below. Taking the material derivative of both sides of Eq. (24) to characterize the error evolution in the continuous limit:
$\frac{D\epsilon_{p}}{Dt}=\frac{D(\rho e)_{real}}{Dt}-\frac{D\left(p\xi+E_{0}\right)}{Dt}$ (25)
Since $\xi$ and $E_{0}$ satisfy advection equations during the evolution step, $D\xi/Dt=0$ and $DE_{0}/Dt=0$. Substituting these yields:
$\frac{D\epsilon_{p}}{Dt}=\frac{D(\rho e)_{real}}{Dt}-\left(\xi\frac{Dp}{Dt}+p\frac{D\xi}{Dt}+{\frac{DE_{0}}{Dt}}\right)=\frac{D(\rho e)_{real}}{Dt}-\xi\frac{Dp}{Dt}$ (26)
From the fundamental thermodynamic relation $Tds=de-(p/\rho^{2})d\rho$, material derivative of energy is given by:
$\frac{D(\rho e)_{real}}{Dt}=\rho T\frac{Ds}{Dt}-(\rho e+p)\frac{\partial u}{\partial x}$ (27)
Applying this to the error evolution equation leads to:
$\frac{D\epsilon_{p}}{Dt}=\rho T\frac{Ds}{Dt}-(\rho e+p)\frac{\partial u}{\partial x}-\xi\frac{Dp}{Dt}$ (28)
Substituting the definition of the thermodynamic coefficient $\xi=(\rho e+p)/(\rho c^{2})$ (Eq. (15)) into the error evolution equation and simplifying, we obtain:
$\frac{D\epsilon_{p}}{Dt}=\rho T\frac{Ds}{Dt}-\xi\left(\frac{Dp}{Dt}+\rho c^{2}\frac{\partial u}{\partial x}\right)$ (29)
Note that the term within the parentheses corresponds to the 1D pressure evolution equation:
$\frac{Dp}{Dt}+\rho c^{2}\frac{\partial u}{\partial x}=\left(\frac{\partial p}{\partial s}\right)_{\rho}\frac{Ds}{Dt}$ (30)
Combining this with Eq. (29), we arrive at the final error evolution equation:
$\frac{D\epsilon_{p}}{Dt}=\left(\rho T-\xi\left(\frac{\partial p}{\partial s}\right)_{\rho}\right)\frac{Ds}{Dt}$ (31)
Based on Eq. (31), we deduce the following error characteristics: In smooth flow regions, since the flow is isentropic ($Ds/Dt=0$), the growth rate of the re-projection error is analytically zero. Numerically, $\epsilon_{p}$ arises solely from truncation errors in the spatial reconstruction, scaling as $O(\Delta x^{k}\Delta t)\sim O(\Delta x^{k+1})$ (where $k$ denotes the reconstruction order). In discontinuous regions, the growth rate of the re-projection error is proportional to the physical entropy production rate $Ds/Dt$. In shock-capturing schemes, entropy production at shocks is governed by numerical viscosity. Consequently, the energy conservation error introduced by this method in discontinuous regions is of order $O(\Delta x)$ (or first-order in time, $O(\Delta t)$).
@@
In summary, the energy conservation error introduced by this method remains consistent with the inherent numerical error of shock-capturing schemes across the entire domain. This analysis demonstrates that, unlike the DF method, the RFQC scheme avoids spatial energy conservation errors arising from flux inconsistency. All thermodynamic inconsistencies are isolated at the end of each time step (the thermodynamic re-projection step). Furthermore, since we solve advection equations for the thermodynamic coefficients, the convergence order of the re-projection error in smooth regions is higher than that of the DF method (which exhibits an $O(\Delta t)$ error due to frozen thermodynamic relations [10]). Additionally, it is worth noting that although the re-projection error term is related to the pressure evolution equation, this does not imply that the proposed method is equivalent to Pressure-Based methods. We strictly follow conservation laws to evolve conserved variables within each time step, which is fundamentally different from directly evolving pressure. The superior conservation properties of the proposed method will be further demonstrated in the numerical examples in the next section.
@@ 4.1 Convergence Test of Energy Conservation Error
With a CFL number of 0.25, we evaluated the re-projection energy conservation error on grids with $N=32,64,128,256,512,1024$. The results are shown in Fig. 3. The profiles of density and pressure for $N=1024$ are displayed in Figs. 3(a) and 3(b). After $t=0.001\text{s}$, the wave has propagated by 0.1 m. The pressure and density profiles remain smooth and stable, demonstrating that the RFQC method effectively suppresses pressure oscillations across the saturation line where thermodynamic properties vary significantly. Fig. 3(c) shows the convergence of the re-projection energy conservation error $|\epsilon_{p}|$ with respect to the grid size. The error exhibits a clear second-order convergence trend, consistent with the theoretical analysis. Moreover, this error does not grow with simulation time, as shown in Fig. 3(d). For $N=1024$, the summation of the relative energy error $\bar{\epsilon_{p}}=\sum|\epsilon_{p}/(\rho e_{t})|$ oscillates slightly around $1.7\times 10^{-7}$ without an increasing trend.
@@ 4.4 Subcritical Liquid Impingement Verification
Left $1\times 10^{6}$ 500 497.3 50
Right $1\times 10^{6}$ 500 497.3 0
@@ 4.5 Shock-Interface Interaction Problem
Middle $(0.5,0.8)$ $2\times 10^{5}$ 10 832.8 0
Right $(0.8,1.0)$ $1\times 10^{6}$ 10 4083 -100
@@ 5 Conclusion
To address the issue of spurious pressure oscillations prevalent in traditional conservative schemes for real fluid calculations, this paper formalizes a generalized strategy involving energy-pressure freezing, flux evolution, pressure reconstruction, and thermodynamic re-projection. This formulation introduces a re-projection error, providing the necessary energy dissipation to suppress pressure oscillations. The success of such methods relies on the appropriate selection and evolution of the linearization coefficients. The Double Flux (DF) method utilizes the isentropic index to construct frozen coefficients and performs re-projection after a double flux evolution. In contrast, the RFQC method employs the Grüneisen coefficient and evolves it within Shyue’s five-equation framework, thereby effectively eliminating the energy flux inconsistency error present in the DF method. Numerical experience suggests that the choice of the internal energy reference point for the Grüneisen parameter influences the algorithm’s numerical characteristics. Therefore, constructing more optimal frozen coefficients will be a focus of future research.
Theoretical analysis demonstrates that the growth rate of the energy conservation error in the RFQC method is proportional to the entropy production rate. This indicates that in smooth flow regions, the error depends solely on the spatial reconstruction order $k$, achieving $k+1$ order accuracy. In discontinuous regions, the error remains first-order, implying that the method does not introduce additional low-order errors beyond those inherent to shock capturing schemes. Numerical experiments further verify that with a first-order space-time discretization scheme, the energy conservation error in smooth regions exhibits second-order accuracy. However, due to the coefficient freezing strategy, the high-order discretization schemes for the advection equations is supposed to differ from those used in traditional two-phase flow methods. Developing high-order discretization schemes for RFQC and extending the method to three-dimensional calculations under realistic conditions will be the primary directions for our future research.

### v2

@@ Abstract
Abstract    A new Real Fluid Quasi-Conservative (RFQC) finite volume method is developed to address the numerical simulation of real fluids involving shock waves in transcritical and phase-change flows. To eliminate the spurious pressure oscillations inherent in fully conservative schemes, we extend the classic quasi-conservative method, originally designed for two-phase flows, to real fluids governed by arbitrary equations of state (EoS). The RFQC method locally linearizes the real fluid EoS at each grid point and time step, constructing and evolving the frozen Grüneisen coefficient $\Gamma$ and the linearization remainder $E_{0}$ via two advection equations. At the end of each time step, the evolved $\Gamma$ and $E_{0}$ are utilized to reconstruct the oscillation-free pressure field, followed by a thermodynamic re-projection applied to the conserved variables. Theoretical analysis demonstrates that, in smooth regions, the energy conservation error introduced by the RFQC method is a second-order small term dominated by the time-step. In discontinuous regions, this error is determined by the entropy increase rate, thereby maintaining consistency with the inherent truncation error of shock-capturing methods. A series of numerical tests verifies that the method can robustly simulate complex flow processes with only minor energy conservation errors, including transcritical flows, phase transitions, and shock-interface interactions. The RFQC method is proven to be both accurate and robust in capturing shock waves and phase transitions.
@@ 3.4 Re-Projection Error Analysis
In the thermodynamic re-projection step, only the total energy is modified. We define the resulting energy conservation error, $\epsilon_{p}=(\rho e_{t})^{n+1}-(\rho e_{t})^{n+1}_{temp}$, as the re-projection error. We now analyze the convergence properties of this error.
The energy $(\rho e_{t})^{n+1}_{temp}$ obtained from the natural evolution of the conservation laws satisfies:
@@
$\epsilon_{p}=\rho^{n+1}e_{real}(\rho^{n+1},p^{n+1})-\left({p^{n+1}(\xi)^{n+1}_{temp}+(E_{0})^{n+1}_{temp}}\right).$ (24)
To provide a discrete analysis of the error, let $\psi(p,s)$ represent the internal energy $(\rho e)_{real}$ as a function of $(p,s)$:
$\psi(p,s)=(\rho e)_{real}(p,s).$ (25)
A Taylor expansion of $\psi$ at the state $(p^{n},s^{n})$ yields:
$\psi^{n+1}=\psi^{n}+\psi_{p}^{n}\Delta p+\psi_{s}^{n}\Delta s+\frac{1}{2}\psi_{pp}^{n}(\Delta p)^{2}+\psi_{ps}^{n}\Delta p\,\Delta s+\frac{1}{2}\psi_{ss}^{n}(\Delta s)^{2}+O(\|\Delta\|^{3}),$ (26)
where the subscripts $p$ and $s$ denote partial derivatives, with $\Delta p=p^{n+1}-p^{n}$ and $\Delta s=s^{n+1}-s^{n}$.
As previously noted, $\xi$ in the real fluid thermodynamic relation is not strictly equal to the Grüneisen parameter. From the isentropic sound speed $c^{2}=(\partial p/\partial\rho)_{s}$ and the thermodynamic relation $(\partial(\rho e)/\partial\rho)_{s}=h$, we have:
$\left(\frac{\partial(\rho e)}{\partial p}\right)_{s}=\frac{(\partial(\rho e)/\partial\rho)_{s}}{(\partial p/\partial\rho)_{s}}=\frac{h}{c^{2}}=\xi=\psi_{p}.$ (27)
This indicates that the linearized internal energy-pressure relation $\rho e=p\xi+E_{0}$ matches both the function value and the first-order pressure derivative along the isentrope, ensuring the correct wave speeds and structures in the Riemann solver.
Considering the advection relation and neglecting discretization higher-order errors, the parameters along a streamline satisfy:
$(\xi)^{n+1}_{temp}=\xi^{n},\quad(E_{0})^{n+1}_{temp}=E_{0}^{n}.$ (28)
Consequently, the second term in Eq. 24 becomes:
$p^{n+1}(\xi)^{n+1}_{temp}+(E_{0})^{n+1}_{temp}$ $=p^{n+1}\xi^{n}+E_{0}^{n}$ (29)
$=(p^{n}\xi^{n}+E_{0}^{n})+\xi^{n}\Delta p$
$=\psi^{n}+\psi_{p}^{n}\Delta p.$
Substituting Eqs. 26 and 29 into the error definition (Eq. 24) yields the local error expansion:
$\epsilon_{p}=\left(\psi_{s}^{n}\Delta s+\frac{1}{2}\psi_{pp}^{n}(\Delta p)^{2}+\psi_{ps}^{n}\Delta p\,\Delta s+\frac{1}{2}\psi_{ss}^{n}(\Delta s)^{2}\right)+O(\|\Delta\|^{3}).$ (30)
Eq. 30 demonstrates that the re-projection error stems primarily from the first-order entropy term $\psi_{s}^{n}\Delta s$ due to non-isentropic deviations and the second-order pressure term $\psi_{pp}^{n}(\Delta p)^{2}$ caused by the nonlinear curvature of the EoS. The error characteristics in different regions are summarized as follows:
  • Isentropic regions: Analytically, $\Delta s=0$, and the error is dominated by the EoS curvature:
    $\epsilon_{p}=\frac{1}{2}\psi_{pp}^{n}(\Delta p)^{2}=\frac{1}{2}\left(\frac{1}{\rho c^{2}}-\frac{\xi}{c^{2}}\left(\frac{\partial c^{2}}{\partial p}\right)_{s}\right)^{n}(\Delta p)^{2}.$ (31)
    In smooth flows, $\Delta p=O(\Delta t)$, resulting in a local error of $O(\Delta t^{2})\sim O(\Delta x^{2})$. The accumulated global error satisfies $\sum\epsilon_{p}=O(\Delta t)$, indicating first-order global conservation behavior.
  • Non-isentropic regions: In the presence of shocks or numerical dissipation, the error is dominated by the first-order entropy change:
    $\epsilon_{p}=\psi_{s}^{n}\Delta s=\left(\rho T-\xi\left(\frac{\partial p}{\partial s}\right)_{\rho}\right)^{n}\Delta s.$ (32)
    Here, the thermodynamic relations $\mathrm{d}(\rho e)=h\,\mathrm{d}\rho+\rho T\,\mathrm{d}s$ and $\mathrm{d}p=c^{2}\mathrm{d}\rho+(\partial p/\partial s)_{\rho}\mathrm{d}s$ are utilized. In shock-capturing schemes, the entropy change is controlled by numerical viscosity; thus, the integrated error remains of the same order as the inherent truncation error of the shock-capturing scheme.
@@ A Taylor expansion of $\psi$ at the state $(p^{n},s^{n})$ yields:
In summary, the thermodynamic conservation error of the proposed RFQC scheme is locally second-order and globally first-order dominated by the time step in smooth isentropic regions. In discontinuous regions, the error growth rate is proportional to the entropy change $\Delta s$, which is consistent with the numerical dissipation required for shock capturing. This analysis indicates that, unlike the DF methods, the present approach avoids spatial energy conservation errors arising from flux inconsistency. All thermodynamic inconsistencies are isolated at the end of each time step within the re-projection step. Furthermore, the RFQC method strictly follows conservation laws to evolve conserved variables within each time step, which fundamentally distinguishes it from non-conservative methods that directly evolve pressure. The superior conservation properties of the proposed method will be further demonstrated in the numerical examples in the next section.
@@ 4.1 Convergence Test of Energy Conservation Error
With a CFL number of 0.25, we evaluated the re-projection energy conservation error on grids with $N=32,64,128,256,512,1024$, for which the time step scales linearly with the grid spacing. The results are shown in Fig. 3. The profiles of density and pressure for $N=1024$ are displayed in Figs. 3(a) and 3(b). After $t=0.001\text{s}$, the wave has propagated by 0.1 m. The pressure and density profiles remain smooth and stable, demonstrating that the RFQC method effectively suppresses pressure oscillations across the saturation line where thermodynamic properties vary significantly. Fig. 3(c) shows the convergence of the re-projection energy conservation error $|\epsilon_{p}|$ with respect to the grid size. The error exhibits a clear second-order convergence trend, consistent with the theoretical analysis. Moreover, this error does not grow with simulation time, as shown in Fig. 3(d). For $N=1024$, the summation of the relative energy error $\bar{\epsilon_{p}}=\sum|\epsilon_{p}/(\rho e_{t})|$ oscillates slightly around $1.7\times 10^{-7}$ without an increasing trend.
@@ 4.4 Subcritical Liquid Impingement Verification
Left $1\times 10^{6}$ 550 497.3 50
Right $1\times 10^{6}$ 550 497.3 0
@@ 4.5 Shock-Interface Interaction Problem
Middle $(0.5,0.8)$ $2\times 10^{5}$ 5 832.8 0
Right $(0.8,1.0)$ $2\times 10^{5}$ 5 832.8 -800
@@ 5 Conclusion
To address the issue of spurious pressure oscillations prevalent in traditional conservative schemes for real fluid calculations, this paper formalizes a generalized strategy involving energy-pressure freezing, flux evolution, pressure reconstruction, and thermodynamic re-projection. This formulation introduces a re-projection error, which acts as an effective mechanism to suppress spurious pressure oscillations. The success of such methods relies on the appropriate selection and evolution of the linearization coefficients. The Double Flux (DF) method utilizes the isentropic index to construct frozen coefficients and performs re-projection after a double flux evolution. In contrast, the RFQC method employs the Grüneisen coefficient and evolves it within Shyue’s five-equation framework, thereby effectively eliminating the energy flux inconsistency error present in the DF method. Numerical experience suggests that the choice of the internal energy reference point for the Grüneisen parameter influences the algorithm’s numerical characteristics. Therefore, constructing more optimal frozen coefficients will be a focus of future research.
Theoretical analysis demonstrates that the energy conservation error of the RFQC method is determined by the entropy production rate and the nonlinear curvature of the real fluid EoS. In smooth flow regions, the local error introduced by the thermodynamic projection is second-order with respect to the time step; in discontinuous regions, the error remains first-order, implying that the conservation error is dominated by time discretization and introduces no additional low-order spatial errors beyond those inherent to shock capturing schemes. Numerical experiments further verify that with a first-order space-time discretization scheme, the energy conservation error in smooth regions exhibits second-order accuracy. However, due to the coefficient freezing strategy, the high-order discretization schemes for the advection equations should differ from those used in traditional two-phase flow methods. Developing high-order discretization schemes for RFQC and extending the method to three-dimensional calculations under realistic conditions will be the primary directions for our future research.